CAVITY RINGDOWN SPECTROSCOPY AND KINETICS OF HO_{2}+HCHO: DETECTION OF THE ν_{1} AND Ã-X̃ BANDS OF HOCH_{2}OO

\underline{MATTHEW K. SPRAGUE}^{a}, MITCHIO OKUMURA, California Institute of Technology, Division of Chemistry, MC 127-72, Pasadena, CA 91125; and STANLEY P. SANDER, Jet Propulsion Laboratory, California Institute of Technology, MS 183-901, Pasadena, CA 91109.

The reactions of HO_{2} with carbonyl compounds are believed to be a sink for carbonyl compounds in the upper troposphere and lower stratosphere. These reactions proceed through a hydrogen bound intermediate before isomerizing.^{b} The reaction of HO_{2} + formaldehyde (HCHO) serves as a prototype for this class of reactions, forming the isomerization product hydroxymethylperoxy (HOCH_{2}OO, HMP). Previous studies measured the spectrum and kinetics of HMP using either FTIR detection of the end products^{c} or direct detection of HMP by the unstructured B̃-X̃ transition.^{d}^{,}^{e} Despite these studies, considerable uncertainty exists in the rate constant of HMP formation (80%, 2σ).^{f} In this talk, we report the first detection of the ν_{1} (OH stretch) and Ã-X̃ electronic spectra of the HMP radical. The OH stretch spectrum is broad and featureless, while the Ã(0)-X̃(0) origin and combination band with the OOCO torsion Ã(n_{OOCO}=1)-X̃(0) are rotationally resolved. Quantum chemistry calculations have been performed on both the Ã and X̃ states as a function of the OOCO and HOCO dihedral angles to estimate the Ã-X̃ transition frequency and to assess the coupling between the two torsional modes. We also present kinetics data showing the rates of production and destruction of HMP.

aSupport from the NDSEG Fellowship, California Air Resources Board Contracts 03-333 and 07-730, and NASA Upper Atmosphere Research Program Grants NAG5-11657, NNG06GD88G and NNX09AE21G are gratefully acknowledged

bI. Hermans, J. F. Muller, T. L. Nguyen, P. A. Jacobs, and J. Peeters. J. Phys. Chem. A 2005, 109, 4303.

cF. Su, J. G. Calvert, and J. H. Shaw J. Phys. Chem. 1979, 83, 3185.

dB. Veyret, R. Lesclaux, M. T. Rayez, J. C. Rayez, R. A. Cox, and G. K. Moortgat J. Phys. Chem. 1989, 93, 2368.

eJ. P. Burrows, G. K. Moortgat, G. S. Tyndall, R. A. Cox, M. E. Jenkin, G. D. Hayman, and B. Veyret J. Phys. Chem. 1989, 93, 2375

fS. P. Sander, B. J. Finlayson-Pitts, D. M. Golden, R. E. Huie, C. E. Kolb, M. J. Kurylo, M. J. Molina, et al. Chemical Kinetics and Photochemical Data for Use in Atmospheric Studies, Evaluation Number 16, Jet Propulsion Laboratory, 2009